# Relative_Constructibility_via_Restricted_ASMs.bib

### v1

@@ 2.5 Input and Output
In Example 2.13, we are able to easily read the input of a Turing machine off its corresponding initial state in the ASM representation. Similarly, we expect to easily read the output off any of its final states. These are self-evident desiderata if we were to port the input-output paradigm of computability to ASMs.
To simplify things, let us require any input or output of an ASM $\mathfrak{M}$ to be a subset of $b(\mathfrak{M})$.
@@
A loading function of an ASM $\mathfrak{M}$ is a function $l$ from $\mathcal{P}(b(\mathfrak{M}))$ onto $I(\mathfrak{M})$, such that for some first-order formula $\phi$ with one free variable $x$ over $\sigma(\mathfrak{M})$,
@@
An unloading function of an ASM $\mathfrak{M}$ is a function $u$ from $S(\mathfrak{M})$ into $\mathcal{P}(b(\mathfrak{M}))$, such that for some first-order formula $\phi$ with one free variable $x$ over $\sigma(\mathfrak{M})$,
@@
Let $\mathfrak{M}$ be an ASM. If $\phi$ is a witness for $l_{1}$ and $l_{2}$ as loading functions of $\mathfrak{M}$, then $l_{1}=l_{2}$.
@@
Let $\mathfrak{M}$ be an ASM. If $\phi$ is a witness for $u_{1}$ and $u_{2}$ as unloading functions of $\mathfrak{M}$, then $u_{1}=u_{2}$.
Let $\mathfrak{M}$ be an ASM, $\phi$ be a witness for a loading function of $\mathfrak{M}$, and $\mathfrak{T}(\mathfrak{M})=(\emptyset,\sigma(\mathfrak{M}),\dot{\mathcal{U}},\vartheta)$. Assume $|b(\mathfrak{M})|>1$.
@@
($\wedge(\neg c_{0}=0_{0}\implies(c_{1}=c_{0}\wedge\forall x\ (\mathrm{In}_{0}(x)\iff\mathrm{In}_{1}(x))\wedge\theta))\urcorner\text{,}$
@@
$\varphi:=\bigwedge\{\ulcorner\forall x\ (\neg R(x))\urcorner:\ulcorner R\urcorner\neq\ulcorner\mathrm{In}\urcorner\text{ and }\ulcorner R\urcorner\text{ is a relation symbol in }\sigma(\mathfrak{M})\}\text{.}$
@@
In accordance to the discussion above, we want an ASM $\mathfrak{M}$ to have a signature $\sigma(\mathfrak{M})$ containing two distinguished unary relation symbols $\ulcorner\mathrm{In}\urcorner$ and $\ulcorner\mathrm{Out}\urcorner$, representing the input and output tapes respectively. Moreover, we want the specification of $\mathfrak{M}$ to include a first-order sentence $\phi^{d}_{\mathfrak{M}}$ over $\sigma(\mathfrak{M})\setminus\{\ulcorner\mathrm{In}\urcorner,\ulcorner\mathrm{Out}\urcorner\}$ that describes the default “non-tape” configuration of $\mathfrak{M}$. As in the case of transition functions, we want the loading function described by $\phi^{d}_{\mathfrak{M}}$ to be absolute for transitive models of $\mathsf{ZFC}$. To do so, we want to first impose a greater amount of structure on the states of an ASM.
@@ 2.6 Constraining the Interpretation Constraints
($\wedge(\neg(\forall x\ (R_{0}(x)\iff\mathrm{Out}_{0}(x)))\implies$
@@
Then basically, $\mathfrak{M}$ is a machine that starts with $\ulcorner R\urcorner$ holding $A$, before copying $A$ onto the output tape. As it does this regardless of input, $\mathfrak{M}$ computes $A$ from any input, even if $A$ is a very complicated set by any other measure.
In the usual input-output paradigm of computability where the output complexity should depend solely on the input complexity, Example 2.24 cautions against hiding too much information in the states of an ASM. One way to fix this is to enforce certain definability constraints on the interpretation constraints.
@@
Justifiably, we shall stipulate that whenever $(A,z)\in ran(\vartheta)$, $z=0$ and $A=\kappa^{n}$ for some $n<\omega$. We are now ready to formally define our modified version of an ASM.
@@
A sequential restricted abstract state machine — or sequential RASM (henceforth just RASM) — $\mathfrak{M}$ is fully specified by a limit ordinal $\kappa(\mathfrak{M})$, a finite signature $\sigma(\mathfrak{M})$, and sentences $\phi^{\tau}_{\mathfrak{M}}$ and $\phi^{d}_{\mathfrak{M}}$ such that
@@
Fix any limit ordinal $\kappa$. Notice that there are only countably many RASMs $\mathfrak{M}$ satisfying $\kappa(\mathfrak{M})=\kappa$. Indeed, this countable set of RASM specifications is absolute for transitive models of $\mathsf{ZFC}$ containing $\kappa$.
@@
If $\mathfrak{M}=(\kappa(\mathfrak{M}),\sigma(\mathfrak{M}),\phi^{\tau}_{\mathfrak{M}},\phi^{d}_{\mathfrak{M}})$ is an RASM, define
@@
[…]
@@
  • $u_{\mathfrak{M}}:=\{&(s,A)\in S(\mathfrak{M})\times\mathcal{P}(\kappa(\mathfrak{M})):\\ &\forall a\in\kappa(\mathfrak{M})\ (a\in A\iff s\models\ulcorner\mathrm{Out}(a)\urcorner)\}\text{.}\end{aligned}$
The next proposition highlights the important properties of RASM, based on what we have discussed so far.
@@
Let $\mathfrak{M}$ be an RASM. Then the following are absolute for transitive models of $\mathsf{ZFC}$ containing $\mathfrak{T}(\mathfrak{M})$:
@@
Now that RASMs have been officially defined, it is imperative to revisit Example 2.13, for any notion of generalised algorithm that cannot represent an arbitrary Turing cannot be too useful.
@@
{($\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=\mathrm{Pre}(h_{0}):$
@@
{($\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=\mathrm{Suc}(h_{0}):$
@@
{($\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=\langle j^{\prime}\rangle:$
@@
{($\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=h_{0}\wedge(\neg x=x):$
@@
{($\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=h_{0}\wedge x=x:$
@@
A rotuine check would reveal that the aforementioned details are sufficient to specify $\mathfrak{M}$ as an RASM, and that they allow $\mathfrak{M}$ to accurately simulate the Turing machine $M$.
From the next subsection onwards, the construction of RASMs used in proofs and arguments can be too cumbersome to formally spell out. In such cases, only high level details will be given. The reader should always try to convince themselves that those details can be properly realised by RASMs.
@@ 2.7 Transfinite Runs
Consider a finite terminating run $\vec{s}=(s_{1},\dots,s_{n})$ of an RASM $\mathfrak{M}$. Then by combining witnesses to $\phi^{\tau}_{\mathfrak{M}}$ being bounded for $\sigma(\mathfrak{M})$, we can conjure witnesses to some sentence $\phi$ being bounded for $\sigma(\mathfrak{M})$, such that using $\phi$ in place of $\phi^{\tau}_{\mathfrak{M}}$ in the definition of $\mathfrak{M}$ gives us an RASM $\mathfrak{M}^{\prime}$ with $\tau_{\mathfrak{M}^{\prime}}(s_{1})=s_{n}$. Therefore, as far as computability is concerned, allowing finite steps of arbitrary sizes in computations is no more powerful than allowing only one-step computations. This goes to show that RASMs are in fact built for runs of infinite lengths.
Given an infinite ordinal $\delta$, what does it mean to have an RASM computation of length $\delta$? Naturally, every state of the computation must depend on the states prior. For successor ordinals $\alpha<\delta$, the $\alpha$-th state of the computation depends solely on its unique previous state, according to the transition function of the RASM. But what about limit ordinals less than $\delta$? How do we describe the states associated with these ordinals in terms of states occurring in the computation before them? This is equivalent to asking how we are going to process the limit of a sequence of states.
Fix an RASM $\mathfrak{M}$, a limit ordinal $\delta$, and let $Sq$ be a sequence of states of $\mathfrak{M}$. We want to have a uniform definition of $lim(Sq)\in S(\mathfrak{M})$. One natural way to do so is to take ordinal limits point-wise (according to the order topology on $\kappa(\mathfrak{M})$ induced by $\in$), while treating relations as boolean functions. The obvious hindrance to this method is, the limit of a sequence of ordinals does not always exist. However, for various purposes, there are suitable ersatzes.
@@
Let $T$ be a function from some ordinal into $ORD$. Then $lim\ inf(T)$, $lim\ sup(T)$ and $lim(T)$ refer respectively to the limit inferior, the limit superior and the limit of $T$ as a sequence, with respect to the order topology induced on $\kappa(\mathfrak{M})$ by $\in$.
@@
If $T$ is a function from some ordinal into $ORD$, then $lim\ inf(T)$ and $lim\ sup(T)$ always exist.
@@
then $lim\ inf(T)$ and $lim\ sup(T)$ always exist in $\kappa(\mathfrak{M})$.
Point-wise $lim\ inf$ and $lim\ sup$ operations are commonly used to define the limit stages of transfinite computations. Hamkins and Lewis, in [ittm], defined each cell in the tape of an infinite time Turing machine to take the $lim\ sup$ of its previous values. Koepke, in his definition of ordinal Turing machines ([koepke1] and [koepke2]), instead require each cell of the machine tape to take the $lim\ inf$ of its previous values. Point-wise $lim\ inf$ and $lim\ sup$ always exist in these generalised Turing machines, for computations shorter than the lengths of their tapes. This is because single-step movements of these machines are always small, and consequently the hypothesis of Fact 2.32 applies.
Hereon, we shall identify relations with boolean functions and constants with 00-ary functions wherever convenient, in part so that only function symbols need to be considered in definitions and case analyses.
@@
Let $\mathfrak{M}$ be an RASM and $Sq$ be a function from some ordinal into $S(\mathfrak{M})$. Given $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ an $n$-ary function symbol, denote by $\ulcorner R^{\ulcorner X\urcorner}\urcorner$ a distinguished $(n+2)$-ary relation symbol and define the relation
@@
Let $\mathfrak{M}$ be an RASM. Given $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ an $n$-ary function symbol, we say $\phi$ is a simple limit formula for $(\ulcorner X\urcorner,\mathfrak{M})$ iff $\phi$ is a formula over $\{\ulcorner\in\urcorner,\ulcorner R^{\ulcorner X\urcorner}\urcorner\}$ with free variables $x_{1},\dots,x_{n+1}$.
@@
for all functions $Sq$ from a limit ordinal $<\kappa(\mathfrak{M})$ into $S(\mathfrak{M})$ and all $(\alpha_{1},\dots,\alpha_{n+1})\in\kappa(\mathfrak{M})^{n+1}$ such that
@@
It is immediate from Definition 2.36 and our assumption that $\sigma(\mathfrak{M})$ contains only function symbols, that a simple limit of a $<\kappa(\mathfrak{M})$-length ordinal sequence of states of $\mathfrak{M}$ must be unique if it exists. Furthermore, for any such sequence $Sq$ and any set $Y$, whether $Y$ witnesses the existence of a simple limit of $Sq$, as well as the identity of said limit if it exists, is absolute for transitive models of $\mathsf{ZFC}$ containing $Sq$.
@@
Let $\mathfrak{M}$ be an RASM. For each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$ that preserves ordinal limits, such that whenever $Sq$ is a function from $\delta<\kappa(\mathfrak{M})$ into $S(\mathfrak{M})$, $\{\phi_{lim}^{\ulcorner X\urcorner}:\ulcorner X\urcorner\in\sigma(\mathfrak{M})\}$ witnesses the existence of a simple limit $s$ of $Sq$ satisfying
@@
By the definition of an RASM and the observation that
$lim\ inf(T)=lim(T)$
whenever $T$ is a function from some ordinal in $ORD$ and $lim(T)$ exists. The reader should check that the $lim\ inf$ values specified by the proposition relative to $\ulcorner X\urcorner$ and $Sq$, in case they exist, are definable over $(\kappa(\mathfrak{M});\in,R_{Sq}^{\ulcorner X\urcorner})$. In fact, there is one such definition that depends only on $\ulcorner X\urcorner$ and not on $\mathfrak{M}$. ∎
Colloquially, Proposition 2.37 tells us that every RASM is strongly closed under limit inferiors. The following definitions generalise this concept.
@@
Let $\mathfrak{M}$ be an RASM. Then $\mathfrak{M}$ is strongly closed under a set $Y$ iff for each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is a simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$, such that
@@
and whenever $Sq$ is a function from some limit ordinal $<\kappa(\mathfrak{M})$ into $S(\mathfrak{M})$, there is $s\in S(\mathfrak{M})$ witnessed by $Y$ to be a simple limit of $Sq$.
@@
Let $\mathfrak{M}$ be an RASM. Then $\mathfrak{M}$ is practically closed under a set $Y$ iff for each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is a simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$, such that
@@
Clearly, for any RASM $\mathfrak{M}$ and any set $Y$,
@@
We are ready to define transfinite terminating runs of RASMs.
@@
Let $\mathfrak{M}$ be an RASM that is practically closed under $Y$. A short terminating run of $(\mathfrak{M},Y)$ with input $A$ and output $B$ is a function $Sq$ from some successor ordinal $\delta<\kappa(\mathfrak{M})$ into $S(\mathfrak{M})$ such that
@@
Suppose $\mathfrak{M}$ is an RASM practically closed under $Y$. Then for some signature $\sigma^{\prime}$, as long as $\mathfrak{M}^{\prime}$ is an RASM satisfying
@@
and $Z$ is a set of simple limit formulas preserving ordinal limits with $\mathfrak{M}^{\prime}$ strongly closed under $Z$, there is an RASM $\mathfrak{M}^{*}$ such that
  1. (1)
@@
Let $\mathfrak{M}^{\prime}$ and $Z$ be as given by the proposition. We need to define $\mathfrak{M}^{*}$ so that (1) to (3) are satisfied. We are thus forced to set

### v2

@@ 2.5 Input and Output
In Example 2.13, we are able to easily read the input of a Turing machine off its corresponding initial state in the SeqA representation. Similarly, we expect to easily read the output off any of its final states. These are self-evident desiderata if we were to port the input-output paradigm of computability to SeqAs.
To simplify things, let us require any input or output of a SeqA $\mathfrak{M}$ to be a subset of $b(\mathfrak{M})$.
@@
A loading function of a SeqA $\mathfrak{M}$ is a function $l$ from $\mathcal{P}(b(\mathfrak{M}))$ onto $I(\mathfrak{M})$, such that for some first-order formula $\phi$ with one free variable $x$ over $\sigma(\mathfrak{M})$,
@@
An unloading function of a SeqA $\mathfrak{M}$ is a function $u$ from $S(\mathfrak{M})$ into $\mathcal{P}(b(\mathfrak{M}))$, such that for some first-order formula $\phi$ with one free variable $x$ over $\sigma(\mathfrak{M})$,
@@
Let $\mathfrak{M}$ be a SeqA. If $\phi$ is a witness for $l_{1}$ and $l_{2}$ as loading functions of $\mathfrak{M}$, then $l_{1}=l_{2}$.
@@
Let $\mathfrak{M}$ be a SeqA. If $\phi$ is a witness for $u_{1}$ and $u_{2}$ as unloading functions of $\mathfrak{M}$, then $u_{1}=u_{2}$.
Let $\mathfrak{M}$ be a SeqA, $\phi$ be a witness for a loading function of $\mathfrak{M}$, and $\mathfrak{T}(\mathfrak{M})=(\emptyset,\sigma(\mathfrak{M}),\dot{\mathcal{U}},\vartheta)$. Assume $|b(\mathfrak{M})|>1$.
@@
$\wedge(\neg c_{0}=0_{0}\implies(c_{1}=c_{0}\wedge\forall x\ (\mathrm{In}_{0}(x)\iff\mathrm{In}_{1}(x))\wedge\theta))\urcorner\text{,}$
@@
[…]
@@
In accordance to the discussion above, we want a SeqA $\mathfrak{M}$ to have a signature $\sigma(\mathfrak{M})$ containing two distinguished unary relation symbols $\ulcorner\mathrm{In}\urcorner$ and $\ulcorner\mathrm{Out}\urcorner$, representing the input and output tapes respectively. Moreover, we want the specification of $\mathfrak{M}$ to include a first-order sentence $\phi^{d}_{\mathfrak{M}}$ over $\sigma(\mathfrak{M})\setminus\{\ulcorner\mathrm{In}\urcorner,\ulcorner\mathrm{Out}\urcorner\}$ that describes the default “non-tape” configuration of $\mathfrak{M}$. As in the case of transition functions, we want the loading function described by $\phi^{d}_{\mathfrak{M}}$ to be absolute for transitive models of $\mathsf{ZFC}$. To do so, we want to first impose a greater amount of structure on the states of a SeqA.
@@ 2.6 Constraining the Interpretation Constraints
Choose a limit ordinal $\kappa$ and $A\subset\kappa$. Let $\mathfrak{M}$ be a SeqA with
$\kappa(\mathfrak{M})=\kappa$
$\sigma(\mathfrak{M})=\{\ulcorner\in\urcorner,\ulcorner\mathrm{In}\urcorner,\ulcorner\mathrm{Out}\urcorner,\ulcorner R\urcorner\}$
and associated TCI
$\mathfrak{T}(\mathfrak{M})=(\emptyset,\sigma(\mathfrak{M}),\dot{\mathcal{U}},\vartheta)\text{,}$
@@
$\wedge(\neg(\forall x\ (R_{0}(x)\iff\mathrm{Out}_{0}(x)))\implies$
@@
Then basically, $\mathfrak{M}$ is an algorithm that starts with $\ulcorner R\urcorner$ holding $A$, before copying $A$ onto the output tape. As it does this regardless of input, $\mathfrak{M}$ computes $A$ from any input, even if $A$ is a very complicated set by any other measure.
In the usual input-output paradigm of computability where the output complexity should depend solely on the input complexity, Example 2.24 cautions against hiding too much information in the states of a SeqA. One way to fix this is to enforce certain definability constraints on the interpretation constraints.
@@
Justifiably, we shall stipulate that whenever $(A,z)\in ran(\vartheta)$, $z=0$ and $A=\kappa^{n}$ for some $n<\omega$. We are now ready to formally define our modified version of a SeqA.
@@
[…]
@@
Fix any limit ordinal $\kappa$. Notice that there are only countably many GSeqAs $\mathfrak{M}$ satisfying $\kappa(\mathfrak{M})=\kappa$. Indeed, this countable set of GSeqAs specifications is absolute for transitive models of $\mathsf{ZFC}$ containing $\kappa$.
@@
If $\mathfrak{M}=(\kappa(\mathfrak{M}),\sigma(\mathfrak{M}),\phi^{\tau}_{\mathfrak{M}},\phi^{d}_{\mathfrak{M}})$ is a GSeqA, define
  1. (1)
@@
  3. (3)
    $\begin{aligned} l_{\mathfrak{M}}:=\{&(A,s)\in\mathcal{P}(\kappa(\mathfrak{M}))\times S(\mathfrak{M}):\\ &\forall a\in\kappa(\mathfrak{M})\ (a\in A\iff s\models\ulcorner\phi^{d}_{\mathfrak{M}}\wedge\forall y\ (\neg\mathrm{Out}(y))\wedge\mathrm{In}(a)\urcorner)\}\end{aligned}$
  4. (4)
@@
  5. (5)
    $\begin{aligned} u_{\mathfrak{M}}:=\{&(s,A)\in S(\mathfrak{M})\times\mathcal{P}(\kappa(\mathfrak{M})):\\ &\forall a\in\kappa(\mathfrak{M})\ (a\in A\iff s\models\ulcorner\mathrm{Out}(a)\urcorner)\}\text{.}\end{aligned}$
The next proposition highlights the important properties of GSeqAs, based on what we have discussed so far.
@@
Let $\mathfrak{M}$ be a GSeqA. Then the following are absolute for transitive models of $\mathsf{ZFC}$ containing $\mathfrak{T}(\mathfrak{M})$:
  1. (1)
@@
Now that GSeqAs have been officially defined, it is imperative to revisit Example 2.13, for any notion of generalised algorithm that cannot represent an arbitrary Turing cannot be too useful.
@@
$\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=\mathrm{Pre}(h_{0}):$
@@
$\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=\mathrm{Suc}(h_{0}):$
@@
$\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=\langle j^{\prime}\rangle:$
@@
$\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=h_{0}\wedge(\neg x=x):$
@@
$\bigvee\{(\neg e_{0}=\langle 0\rangle)\wedge q(\mathrm{Out}_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle\wedge x=h_{0}\wedge x=x:$
@@
A routine check would reveal that the aforementioned details are sufficient to specify $\mathfrak{M}$ as a GSeqA, and that they allow $\mathfrak{M}$ to accurately simulate the Turing machine $M$.
From the next subsection onwards, the construction of GSeqAs used in proofs and arguments can be too cumbersome to formally spell out. In such cases, only high-level details will be given. The reader should always try to convince themselves that those details can be properly realised by GSeqAs.
@@ 2.7 Transfinite Runs
Consider a finite terminating run $\vec{s}=(s_{1},\dots,s_{n})$ of a GSeqA $\mathfrak{M}$. Then by combining witnesses to $\phi^{\tau}_{\mathfrak{M}}$ being bounded for $\sigma(\mathfrak{M})$, we can conjure witnesses to some sentence $\phi$ being bounded for $\sigma(\mathfrak{M})$, such that using $\phi$ in place of $\phi^{\tau}_{\mathfrak{M}}$ in the definition of $\mathfrak{M}$ gives us a GSeqA $\mathfrak{M}^{\prime}$ with $\tau_{\mathfrak{M}^{\prime}}(s_{1})=s_{n}$. Therefore, as far as computability is concerned, allowing finite steps of arbitrary sizes in computations is no more powerful than allowing only one-step computations. This goes to show that GSeqAs are in fact built for runs of infinite lengths.
Given an infinite ordinal $\delta$, what does it mean to have a GSeqA computation of length $\delta$? Naturally, every state of the computation must depend on the states prior. For successor ordinals $\alpha<\delta$, the $\alpha$-th state of the computation depends solely on its unique previous state, according to the transition function of the GSeqA. But what about limit ordinals less than $\delta$? How do we describe the states associated with these ordinals in terms of states occurring in the computation before them? This is equivalent to asking how we are going to process the limit of a sequence of states.
Fix a GSeqA $\mathfrak{M}$, a limit ordinal $\delta$, and let $Sq$ be a sequence of states of $\mathfrak{M}$. We want to have a uniform definition of $lim(Sq)\in S(\mathfrak{M})$. One natural way to do so is to take ordinal limits point-wise (according to the order topology on $\kappa(\mathfrak{M})$ induced by $\in$), while treating relations as boolean functions. The obvious hindrance to this method is, the limit of a sequence of ordinals does not always exist. However, for various purposes, there are suitable ersatzes.
@@
Let $T$ be a function from some ordinal into $ORD$. Then liminf(T)lim\ inf(T), limsup(T)lim\ sup(T) and $lim(T)$ refer respectively to the limit inferior, the limit superior and the limit of $T$ as a sequence, with respect to the order topology induced on $\kappa(\mathfrak{M})$ by $\in$.
@@
If $T$ is a function from some ordinal into $ORD$, then liminf(T)lim\ inf(T) and limsup(T)lim\ sup(T) always exist.
@@
then liminf(T)lim\ inf(T) and limsup(T)lim\ sup(T) always exist in $\kappa(\mathfrak{M})$.
Point-wise liminflim\ inf and limsuplim\ sup operations are commonly used to define the limit stages of transfinite computations. Hamkins and Lewis, in [8], defined each cell in the tape of an infinite time Turing machine to take the limsuplim\ sup of its previous values. Koepke, in his definition of ordinal Turing machines ([13] and [14]), instead require each cell of the machine tape to take the liminflim\ inf of its previous values. Point-wise liminflim\ inf and limsuplim\ sup always exist in these generalised Turing machines, for computations shorter than the lengths of their tapes. This is because single-step movements of these machines are always small, and consequently the hypothesis of Fact 2.32 applies.
Hereon, we shall identify relations with boolean functions and constants with 0-ary functions wherever convenient, in part so that only function symbols need to be considered in definitions and case analyses.
@@
Let $\mathfrak{M}$ be a GSeqA and $Sq$ be a function from some ordinal into $S(\mathfrak{M})$. Given $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ an $n$-ary function symbol, denote by $\ulcorner R^{\ulcorner X\urcorner}\urcorner$ a distinguished $(n+2)$-ary relation symbol and define the relation
@@
Let $\mathfrak{M}$ be a GSeqA. Given $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ an $n$-ary function symbol, we say $\phi$ is a simple limit formula for $(\ulcorner X\urcorner,\mathfrak{M})$ iff $\phi$ is a formula over $\{\ulcorner\in\urcorner,\ulcorner R^{\ulcorner X\urcorner}\urcorner\}$ with free variables $x_{1},\dots,x_{n+1}$.
@@
for all functions $Sq$ from a limit ordinal $\kappa(\mathfrak{M})$ into $S(\mathfrak{M})$ and all $(\alpha_{1},\dots,\alpha_{n+1})\in\kappa(\mathfrak{M})^{n+1}$ such that
@@
It is immediate from Definition 2.36 and our assumption that $\sigma(\mathfrak{M})$ contains only function symbols, that a simple limit of a $\kappa(\mathfrak{M})-$length ordinal sequence of states of $\mathfrak{M}$ must be unique if it exists. Furthermore, for any such sequence $Sq$ and any set $Y$, whether $Y$ witnesses the existence of a simple limit of $Sq$, as well as the identity of said limit if it exists, is absolute for transitive models of $\mathsf{ZFC}$ containing $Sq$.
@@
Let $\mathfrak{M}$ be a GSeqA. For each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$ that preserves ordinal limits, such that whenever $Sq$ is a function from $\delta<\kappa(\mathfrak{M})$ into $S(\mathfrak{M})$, $\{\phi_{lim}^{\ulcorner X\urcorner}:\ulcorner X\urcorner\in\sigma(\mathfrak{M})\}$ witnesses the existence of a simple limit $s$ of $Sq$ satisfying
@@
By the definition of a GSeqA and the observation that
liminf(T)=lim(T)lim\ inf(T)=lim(T)
whenever $T$ is a function from some ordinal in $ORD$ and $lim(T)$ exists. The reader should check that the liminflim\ inf values specified by the proposition relative to $\ulcorner X\urcorner$ and $Sq$, in case they exist, are definable over $(\kappa(\mathfrak{M});\in,R_{Sq}^{\ulcorner X\urcorner})$. In fact, there is one such definition that depends only on $\ulcorner X\urcorner$ and not on $\mathfrak{M}$. ∎
Colloquially, Proposition 2.37 tells us that every GSeqA is strongly closed under limit inferiors. The following definitions generalise this concept.
@@
Let $\mathfrak{M}$ be a GSeqA. Then $\mathfrak{M}$ is strongly closed under a set $Y$ iff for each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is a simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$, such that
@@
and whenever $Sq$ is a function from some limit ordinal $\kappa(\mathfrak{M})$ into $S(\mathfrak{M})$, there is $s\in S(\mathfrak{M})$ witnessed by $Y$ to be a simple limit of $Sq$.
@@
Let $\mathfrak{M}$ be a GSeqA. Then $\mathfrak{M}$ is practically closed under a set $Y$ iff for each $n$-ary function symbol $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ there is a simple limit formula $\phi_{lim}^{\ulcorner X\urcorner}$ for $(\ulcorner X\urcorner,\mathfrak{M})$, such that
@@
Clearly, for any GSeqA $\mathfrak{M}$ and any set $Y$,
@@
We are ready to define transfinite terminating runs of GSeqAs.
@@
Let $\mathfrak{M}$ be a GSeqA that is practically closed under $Y$. A short terminating run of $(\mathfrak{M},Y)$ with input $A$ and output $B$ is a function $Sq$ from some successor ordinal $\delta<\kappa(\mathfrak{M})$ into $S(\mathfrak{M})$ such that
  1. (a)
@@
Suppose $\mathfrak{M}$ is a GSeqA practically closed under $Y$. Then for some signature $\sigma^{\prime}$, as long as $\mathfrak{M}^{\prime}$ is a GSeqA satisfying
@@
and $Z$ is a set of simple limit formulas preserving ordinal limits with $\mathfrak{M}^{\prime}$ strongly closed under $Z$, there is a GSeqA $\mathfrak{M}^{*}$ such that
  1. (a)
@@
Let $\mathfrak{M}^{\prime}$ and $Z$ be as given by the proposition. We need to define $\mathfrak{M}^{*}$ so that (a) to (c) are satisfied. We are thus forced to set